OpenStudy (anonymous):

( 2 2/5 ) divided by ( 1 3/4 ) = Please explain.

OpenStudy (anonymous):

1.37142857

OpenStudy (anonymous):

and the fraction would be????

OpenStudy (anonymous):

2 2/5 = 12/5 1 3/4 = 7/4 Dividing the two you get: 12/5 * 4/7 = 48/35

OpenStudy (anonymous):

1 13/35 :) or 48/35 :)